LASER SPECTROSCOPY, AB INITIO CALCULATIONS AND VIBRATIONAL ASSIGNMENTS OF SOME DERIVATIVES OF 2-OXAZOLIN-5-ONES

\underline{V. B. SINGH}, Department of Physics, Udai Pratap Autonomous College, Varanasi 221002; A. K. RAI, S. B. RAI, Department of Physics, Banaras Hindu University, Varanasi-221005; K. N. SINGH, Department of Chemistry, Banaras Hindu University, Varanasi-221005.

The spectroscopic properties of some derivatives of 2-oxazolin-5-ones have been investigated by a combination of experimental spectroscopic methods and ab initio calculations. Vibrational assignments for the ground state of some of these organic molecules have been carried out using Raman spectra and ab initio calculations. A splitting in the carbonyl group (C=O stretching) frequency observed in the vibrational spectra, is explained as Fermi-resonance. Conformational studies regarding the effect of moving the methoxy group in the 2-phenyl-4-(4-methoxy benzylidene)-2-oxazolin-5- one molecule to a different position on the ring has been carried out. The strong fluorescence and the observation of a circular ring around the incident laser radiation indicate some probable stimulated process in the above organic derivative. Laser spectroscopic study of some of these molecules is presently in progress.